1. Water enters a 7-cm-diameter pipe steadily with a uniform velocity of 2 m/s and exits with the turbulent flow velocity distribution given by u = umax (I - r/R)1/7. If the pressure drop along the pipe is 10 kPa. determine the drag force exerted on the pipe by water flow.



2. A 5-cm-diameter horizontal jet of water with a velocity of 30 m/s relative to the ground strikes a flat plate that is moving in the same direction as the jet at a velocity of 20 m/s. The water splatters in all directions in the plane of the plate. How much force does the water steam exert on the plate?


1
Expert's answer
2021-12-20T10:26:24-0500

1.

"F=\rho Av^2\frac u p=912 N."

2.

"F=\rho \pi d^2(v-v_0)^2=785~N."